OpenStudy (thatonegirl_):

1.55 billion written out??

jimthompson5910 (jim_thompson5910):

@Outh-1 you wrote out 155 billion

OpenStudy (anonymous):

OH

OpenStudy (thatonegirl_):

are you sure it isn't 1,550,000,000

OpenStudy (anonymous):

IM DUMB DONT LISTEN TO ME

OpenStudy (thatonegirl_):

or is it like 1,500,500,000 loll

jimthompson5910 (jim_thompson5910):

1.55 billion = 1.5 * 10^9 = 1,550,000,000

OpenStudy (thatonegirl_):

Okay thank you :)

jimthompson5910 (jim_thompson5910):

oops typo 1.55 billion = 1.55 * 10^9 = 1,550,000,000

jimthompson5910 (jim_thompson5910):

np